# A Statistician’s Overview of Mechanistic-Informed Modeling

### v1

@@ 5.1 Data Models
Assume we have up to $p$ different datasets at time $t$ given by $\{\bm{z}_{1t},\ldots,\bm{z}_{pt}\}$ that are all thought to be informative about our latent process of interest, $\bm{u}_{t}$. Here, we assume these datasets are conditionally independent given the latent spatio-temporal process. Specifically, $\bm{z}_{jt}|\bm{u}_{t},\bm{\theta}_{zj}\sim\;\;\mathcal{D}(h_{j}(\bm{u}_{t});\bm{\theta}_{zj}),$ where for $j=1,\ldots,p$, $\mathcal{D}(h_{j}(\bm{u}_{t});\bm{\theta}_{zj})$ is some data distribution (e.g., Gaussian, Poisson, log-normal, etc.) that is conditioned on the latent spatio-temporal process, $\bm{u}_{t}$, and parameters $\bm{\theta}_{zj}$, and where $h_{j}(\;)$ is a mapping function between the latent process and the data that can account for missing observations, nonlinear transformations, and change-of-support (spatial resolution) (for more background see Cressie & Wikle,, 2011).
The important aspect of these data models is they can accommodate multiple data sources (data fusion) of different types (e.g., count data, positive support data, Gaussian, e.g., Bradley, (2022)), different supports (spatial resolution and alignment, e.g., Wikle & Berliner, (2005)), measurement errors, and missing data. In addition, as mentioned previously, one can include output from a mechanistic model (e.g., an earth-system model) as “data”, which is another way to incorporate mechanistic information into the analysis (e.g., see Wikle et al.,, 2001; Wikle & Berliner,, 2005; Fuentes & Raftery,, 2005; Leeds et al.,, 2014, for examples). Similarly, if one assumes the latent dyanmic process in Section 5.2.2 is governed by a NN, then as in the case of B-PINNs, we have additional data models that condition on the neural network parameters as discussed in Section 4.2.
@@ 5.2 Process Models
We let the “process” part of the model be decomposed into three terms; an overall time-varying spatial mean, a latent dynamic process, and a residual (discrepancy) process, respectively as:
$\psi(\bm{u}_{t})=\bm{\mu}_{t}+\bm{\Phi}{\bm{u}}_{d,t}+\bm{\nu}_{t},$ (13)
where $\psi(\;)$ is a link function as in generalized linear mixed models (e.g., identity, logit, log, etc.) or an activation function as in neural modeling (e.g., sigmoid, tanh, ReLU, etc.), $\bm{u}_{d,t}$ is the latent dynamic process, and $\bm{\nu}_{t}$ is the discrepancy process. In this case, $\bm{\Phi}$ is a $n\times n_{d}$ matrix that can represent $n_{d}$ spatial basis functions, the identity (in which $n_{d}=n$), or some other linear mapping between the $\bm{u}_{d,t}$ dynamic process and the transformed process. Each of these terms can be specified in multiple ways as described below. Here, we assume $\bm{u}_{t}$ is an $n\times 1$ vector composed of the latent $u$-process at $n$ spatial locations (but as noted in Section 2.1, this can be a multivariate latent vector).
@@ 5.2.1 Mean
The mean process $\bm{\mu}_{t}$ can accommodate covariates or other “fixed effects”; e.g., $\bm{\mu}_{t}=\bm{X}_{t}\bm{\beta}_{t}$, where $\bm{X}_{t}$ is an $n\times q$ covariate matrix with potentially time-varying covariates, and $\bm{\beta}_{t}$ is a $q$-dimensional vector (that may be time varying). In this case, we would specify a prior distribution $[\bm{\beta}_{t}]$ (which depends on the problem). We use the term “covariates” quite loosely here, and note that $\bm{X}_{t}$ this term can accommodate seasonality, temporal trends, one hot encoding, exogenous covariate effects, etc. In some cases, one may not need this term in the model, or it might not depend on time (e.g., it might represent a simple constant offset).
5.2.2 Latent Dynamic Process
The latent dynamic component of the model ${\bm{u}}_{d,t}$ is the most important as it represents the underlying spatio-temporal dynamic mechanism driving the process. This is the main place where we impart mechanistic information (e.g., a PINN or a mechanistically-motivated spatio-temporal statistics model). Note the $n\times n_{d}$ matrix $\bm{\Phi}$ in (13) can represent a mapping matrix or, as is common in statistical DSTMs, a matrix of spatial basis functions (typically, low-rank). In that case, the $n_{d}$-dimensional dynamical process ${\bm{u}}_{d,t}$ operates in some basis space (e.g., see the overview in Cressie & Wikle,, 2011).
The dynamical process can be a PINN as described in Section 4.1 or B-PINN as in Section 4.2, in which case we write with a slight abuse of notation,
${\bm{u}}_{d,t}={\bm{u}}_{NN,t}(\bm{\theta}_{W}),$ (14)
where we are not explicitly writing out the initial and boundary conditions, but those can be incorporated in the PINN as described previously. In this case, the PDE constraint is incorporated into “observations” of the PDE residuals (8) as before (e.g., we can include a data model for these residuals and specify the observations to be zero unless there is a known PDE forcing term, or specify a prior on the time derivative of the PDE). This then accommodates the PDE dynamical structure and we note there are additional $\bm{\theta}_{m}$ parameters that come in the residual process from the PDE. As with the B-PINN framework, the residual variance term ($\sigma^{2}_{r}$) allows one to control the degree of determinism in the PDE representation. In full generality it should be emphasized that the latent dynamic process need not be deterministic. For example, as in Section 3.2.1 consider a mechanistically-motivated DSTM, which we write here as
${\bm{u}}_{d,t}={\cal M}({\bm{u}}_{d,t-1},B({\bm{u}}_{d,t};\bm{\theta}_{b});\bm{\theta}_{m})+\bm{\eta}_{t},$ (15)
where ${\cal M}(\;)$ represents a linear or nonlinear operator that depends on parameters $\bm{\theta}_{m}$ and boundary conditions $B({\bm{u}}_{d,t};\bm{\theta}_{b})$ (that can be learned), and the relationship is explicitly Markovian with a zero-mean Gaussian error process $\bm{\eta}_{t}$ with covariance matrix (potentially time-varying) $\bm{Q}_{t}(\bm{\theta}_{q})$ that depends on parameters $\bm{\theta}_{q}$ (e.g., for an overview see Cressie & Wikle,, 2011; Wikle et al.,, 2019). We also must specify explicitly an initial condition for the the process, say ${\bm{u}}_{d,0}\sim\text{Gau}(\bm{a},\bm{\Sigma}_{0})$ as is usual with spatio-temporal dynamic models.
An important difference between the PINN formulation in (14) and the mechanistically-motivated DSTM in (15) is that the NN model considers all times jointly (i.e., treating time as exchangeable), whereas the dynamic model is Markovian and explicitly considers time conditionally.
@@ 5.2.3 Discrepancy Process
The discrepancy process term $\bm{\nu}_{t}$ is important because it respects the fact that our latent dynamical model represented by the process ${\bm{u}}_{d,t}$ surely does not completely account for all of the latent spatio-temporal dependence in real-world data (i.e., “all models are wrong”). This term could be accommodated by another NN, say $\bm{\nu}_{t}=\bm{\delta}_{t}(\cal{S};\bm{\theta}_{\delta})$ as in Zou et al., (2024). Alternatively, it could be a time-varying stochastic process, e.g., $\bm{\nu}_{t}=f_{\nu}(\bm{\nu}_{t-1};\bm{\theta}_{\nu})$ such as the residual autoregresssive process to accommodate small scale dynamics in Wikle et al., (2001), or simply a time-independent spatial GP (e.g., $\bm{\nu}_{t}\sim\text{Gau}(\bm{0},\bm{\Sigma}))$. More generally, this term could be represented by any spatio-temporal random effect distribution, say $\bm{\nu}_{t}\sim[\bm{\theta}_{\nu}]$.
@@ 5.3 Parameter Models
Depending on the models selected in the data and process stages, one can have a variety of parameters that require prior distributions, e.g., $[\bm{\beta}_{t}|\bm{\theta}_{\beta}][\bm{\theta}_{w}][\bm{\theta}_{\nu}][\bm{\theta}_{\delta}],etc.$ In most cases, the parameters closer to the data and process should have informative priors, either because one has scientific knowledge (e.g., a diffusion parameter should change with space and/or time) or because they might be weakly identified and require empirical Bayes or meta-learning priors (e.g., as in the case of PINNs). Typically, we require regularization for these priors as there are likely to be quite a lot of them, or as mentioned in the context of physical-statistical models, we might have mechanistic information for processes (e.g., parameters must obey a turbulent scaling relationship as in Wikle et al., (2001)).
Deep BHMs as described here have multiple hyperparameters. These correspond to specifications of the parameters for the priors in the previous stages as well as model architecture (e.g., number of hidden units, levels, etc.) in the case of NN process models. Although in some cases these can be given vague proper prior distributions, they are typically fixed (i.e., point mass priors) and the sensitivity to the choice is examined. These choices are very problem specific.
@@ 5.4 Implementation Considerations
When implementing hybrid neural-statistical models such as GMID-DSTM with a NN dynamic process, multiple factors need to be considered. As with many mathematical and statistical models, model identifiability must be be addressed, specifically with regard to the NN and discrepancy term. For example, in (13), if $\bm{u}_{d,t}$ is parameterized using (14) (with a mechanistic component) and $\bm{\nu}_{t}$ is parameterized using a GP, there is potential for confounding between the two processes. There are multiple ways to mediate this issue. The most natural is to restrict the spatial (temporal, or spatio-temporal) range of the GP to operate on a different scale than the dynamics imposed on $\bm{u}_{d,t}$. This can be accomplished by an informative prior distribution on the argument(s) of the kernel function for the GP.
@@
The next considerations are: how best to incorporate a NN within a statistical model, how to specify the architecture of the NN, and how to select prior distributions on the NN weights (and biases). As discussed in Section 4.2, a Bayesian NN (and by extension the B-PINN) is the most natural method to incorporate the NN. However, all the same challenges facing Bayesian NNs and B-PINNs are exhibited in the hybrid models. Regarding the NN architecture, this is a relatively unexplored area, though Toscano et al., (2025) reviews various neural network architectures that have been considered in the context of PINNs. Within the B-PINN literature, simple NNs, such as narrow 3-layer feed-forward NNs, have been used. This could be an artifact of the bias-variance trade-off, where wide and deep NNs are over-parameterized and have the potential to overfit. As such, empirical results suggest narrow and shallow feed-forward NNs result in the best performance, although this may not be the case for complex dynamics. Another potential reason for using relatively simple NNs in B-PINNs is they are more computationally tractable to fit, and we next discuss computational considerations.
The primary computational tool that has been used for posterior inference in Bayesian hierarchical models is MCMC, such as Gibbs sampling (and variants like Metropolis-within-Gibbs sampling) (Brooks et al.,, 2011). Modern gradient-based methods, such as Hamiltonian Monte Carlo (HMC; Neal et al.,, 2011) and extensions like the No-U-Turn Sampler (Hoffman & Gelman,, 2014), have become prevalent in general for Bayesian inference. Often it is the case that the computational cost of an expensive simulator makes MCMC prohibitively expensive, but an emulator or surrogate model can speed up the computations (Higdon et al.,, 2008; Hooten et al.,, 2011; Gopalan & Wikle,, 2022), especially in physical-statistical models. Moreover, physical-statistical models typically involve GPs, which can be computationally difficult due to cubic complexity of inversion and determinant calculations; however, there are ways to reduce this burden, such as the use of sparse precision matrices, reduced-rank basis representations, nearest neighbor approaches, and Vecchia approximations (e.g., see the review in Heaton et al.,, 2019).
@@
In the B-PINN literature (Yang et al.,, 2021), the two primary computational tools are HMC, which has already been introduced, and variational inference (Blei et al.,, 2017). In the case of variational Bayes, the mean-field approximation (i.e., product of independent Gaussians) for posterior neural network weights might not be realistic in the sense that correlations (which are likely to exist in neural network parameters) are neglected, although variational GPs (Tran et al.,, 2015) can help alleviate this issue. Yang et al., (2021) found gradient-based MCMC methods result in the most optimal performance for B-PINNs and for our example presented below in Section 6, we proceed with this approach.
6 Simulation Experiment: GMID-DSTM
@@
To demonstrate the GMID-DSTM framework presented in Section 5 with a B-PINN dynamic process, we consider a simulation experiment. In particular, we consider a mechanistic dynamic process described by a one-dimensional (in space) Burgers’ equation (Bateman,, 1915; Burgers,, 1948), which is a classic nonlinear PDE given by
$\frac{\partial\tilde{u}(s,t)}{\partial t}+\tilde{u}(s,t)\frac{\partial\tilde{u}(s,t)}{\partial s}-\lambda\frac{\partial^{2}\tilde{u}(s,t)}{\partial s^{2}}=0,$ (16)
@@
where $\lambda$ is the diffusion coefficient. The Burgers’ equation is often used to illustrate dynamics because it is a very simple nonlinear advection-diffusion representation of a fluid dynamics problem that can model features like nonlinear wave propagation and shock formation while remaining relatively easy to analyze and solve. Here, we also include a fixed-effects mean term in the model and a Gaussian process discrepancy to reflect a situation where the true process is not just a known mechanistic model. We also add measurement error to the simulation process. The details of the simulation are given below.
@@
Figure 1: Simulated process $u(s,t)$ (left) and components over one-dimensional space and time: Burgers’ equation $\tilde{u}(s,t)$ (second from left), mean process $\mu(s,t)$ (second from right), and discrepancy process $\nu(s,t)$ (right).
@@
    Latent Dynamic Process ($\tilde{u}(s,t)$): For each spatial and temporal location, we generated the PDE solution to Burgers’ equation (16) numerically (with a spectral solver), denoted by: $\tilde{u}(s,t)$. This numerical solution used initial condition $\tilde{u}(s,0)=\exp(-s^{2})$ and boundary conditions $\tilde{u}(-\pi,t)=\tilde{u}(\pi,t)=0$.
@@
    Discrepancy Process ($\nu(s,t)$): For each time point a spatial discrepancy field was drawn as an independent, zero-mean Gaussian process, i.e., $\nu(s,t)\sim\text{GP}({0},k(s,s^{\prime}))$, where $k(s,s^{\prime})=\sigma^{2}_{\nu}\exp\bigg{(}-\frac{(s-s^{\prime})^{2}}{2\ell_{\nu}^{2}}\bigg{)}$ is a squared-exponential covariance function. Specifically, we select $\sigma^{2}_{\nu}=0.05$, and $\ell_{\nu}=0.15$.
@@
    Simulated Process ($u(s,t)$): The simulated process at each spatial and temporal coordinate is then given as the sum of the mean, latent dynamic process, and discrepancy: $u(s,t):=\bm{x}(s,t)^{\top}\bm{\beta}+\tilde{u}(s,t)+\nu(s,t)$.
@@
    Simulated Observations ($z(s,t)$): We assumed that 50% of the space coordinates are missing (sampled at random spatially and fixed across all time points - which is generally a realistic scenario for simulating environmental processes compared to data being missing completely at random in space and time). We then added mean zero independent, identically distributed Gaussian measurement error to the process values, $u(s,t)$ for the randomly selected spatial and temporal coordinates with measurement error variance $\sigma^{2}_{d}=0.2^{2}$, to get observations $z(s,t)$.
Figure 1 shows the simulated space-time fields of the true process, latent dynamic process, mean, and discrepancy fields with the simulated data and data with missing observations given in Figure 2.
@@
We consider the following BHM to fit these simulated data. First, the data and process models are given by:
[…]
The latent dynamic and discrepancy processes are then given by:
[…]
@@
$\quad\frac{\partial{u}_{NN}(s,t)}{\partial t}\sim\text{Gau}(u_{NN}(s,t)\frac{\partial{u}_{NN}(s,t)}{\partial s}-\lambda\frac{\partial^{2}{u}_{NN}(s,t)}{\partial s^{2}},\sigma^{2}_{r})$
[…]
$\quad k(s,s^{\prime})=\sigma^{2}_{\nu}\exp\bigg{(}-\frac{(s-s^{\prime})^{2}}{2\ell_{\nu}^{2}}\bigg{)},$
@@
In this case, $\bm{H}_{t}$ is an incidence matrix that accommodates missing observations (e.g., see the discussion in Wikle et al.,, 2019). In addition, the following prior distributions are specified:
$\sigma^{2}_{d}$ $\sim\text{Cauchy}(\mu_{d},\gamma_{d})\cdot\text{I}(\ell_{d}<\sigma^{2}_{d}<u_{d}),\qquad$ $\bm{\beta}\sim\text{Gau}(\bm{\mu}_{\bm{\beta}},c_{\bm{\beta}}\bm{I})$
$\lambda$ $\sim\text{Gau}(\mu_{\lambda},\sigma^{2}_{\lambda})\cdot\text{I}(0\leq\lambda<\infty),\qquad$ $\sigma^{2}_{\nu}\sim\text{Cauchy}(\mu_{\nu},\gamma_{\nu})\cdot\text{I}(\ell_{\nu}<\sigma^{2}_{\nu}<u_{\nu})$
$\ell_{\nu}$ $\sim\text{Gau}(\mu_{\ell},\sigma^{2}_{\ell})\cdot\text{I}(\ell_{\ell}\leq\ \ell_{\nu}<u_{\ell}),\qquad$ $\bm{\theta}_{W}\sim\text{Gau}(\bm{\mu}_{W},c_{W}\bm{I}),$

### v2

@@ 5.1 Data Models
Assume we have up to $p$ different datasets at time $t$ given by $\{\bm{z}_{1t},\ldots,\bm{z}_{pt}\}$ that are all thought to be informative about our latent process of interest, $\bm{u}_{t}$. As is common in spatio-temporal statistics, we assume these datasets are conditionally independent given the latent spatio-temporal process. Specifically, $\bm{z}_{jt}|\bm{u}_{t},\bm{\theta}_{zj}\sim\;\;\mathcal{D}(h_{j}(\bm{u}_{t});\bm{\theta}_{zj}),$ where for $j=1,\ldots,p$, $\mathcal{D}(h_{j}(\bm{u}_{t});\bm{\theta}_{zj})$ is some data distribution (e.g., normal, Poisson, log-normal, etc.) that is conditioned on the latent spatio-temporal process, $\bm{u}_{t}$, and parameters $\bm{\theta}_{zj}$. Here, $h_{j}(\;)$ is a mapping function between the latent process and the data that can account for missing observations, nonlinear transformations, and change-of-support (spatial resolution) (for more background see cressie2011statistics). For instance, in the example of Section 6, the function $h(.)$ implicitly is a composition of an incidence matrix (to account for missing data locations) and exponentiation (to provide a link function for the Poisson distribution, which requires a positive rate parameter); in other words, this mapping can accommodate a link function in a generalized linear model context. As discussed in the context of physical-statistical models in Section 3.2.1, data models can accommodate multiple observational data sources as well as mechanistic model output. In addition, as discussed in the context of B-PINNs in Section 4.2, these data models can accommodate “pseudo-data” for PDE residuals, boundary conditions, and initial conditions.
@@ 5.2 Process Models
We let the latent process part of the model be decomposed into three terms; an overall time-varying spatial mean, a PINN dynamic process, and a residual (discrepancy) process, respectively as
$\bm{u}_{t}=\bm{\mu}_{t}+{\bm{v}}_{NN,t}+\bm{\delta}_{t},$ (13)
where we are interested in modeling the latent process at a fixed set of $n$ locations given by the $n\times 1$ vector ${\bm{u}}_{t}$. Then, $\bm{\mu}_{t}$ is the $n$-dimensional time-varying spatial mean, $\bm{v}_{NN,t}$ is the PINN at the $n$ locations of interest at time $t$, and $\bm{\delta}_{t}$ is the discrepancy term. Each of these terms can be specified in multiple ways as described below.
@@ 5.2.1 Mean
The process mean $\bm{\mu}_{t}$ can accommodate covariates or other “fixed effects”; e.g., $\bm{\mu}_{t}=\bm{X}_{t}\bm{\beta}_{t}$, where $\bm{X}_{t}$ is an $n\times q$ covariate matrix with potentially time-varying covariates, and $\bm{\beta}_{t}$ is a $q$-dimensional vector (that may be time varying). In this case, we would specify a prior distribution $\bm{\beta}_{t}]$ (which depends on the application; e.g., see the simulation example in the next section). We use the term “covariates” quite loosely here, and note that $\bm{X}_{t}$ can accommodate seasonality, temporal trends, one hot encoding, exogenous covariate effects, etc. In some cases, one may not need this term in the model, or it might not depend on time (e.g., it could represent a simple constant offset).
5.2.2 PINN Dynamic Process
The dynamic component of the model is the most important as it represents the underlying spatio-temporal mechanism driving the process. This is the main place where we impart mechanistic information via the PINN as described in Section 4.1 or B-PINN as in Section 4.2, denoted as ${\bm{v}}_{NN,t}$. Note that we are not explicitly writing out the initial and boundary conditions, but those can be incorporated in the PINN as described previously. Recall, the PDE constraint can be incorporated into “observations” of the PDE residuals (8) as before (e.g., we can include a data model for these residuals and specify the observations to be zero unless there is a known PDE forcing term) or we can specify a prior on the time derivative of the PDE. This then accommodates the PDE dynamical structure and we note there are additional $\bm{\theta}_{m}$ parameters that come in the residual process from the PDE. As with the B-PINN framework, the residual variance term ($\sigma^{2}_{r}$) allows one to control the degree to which the neural network is constrained to follow the underlying PDE.
@@ 5.2.3 Discrepancy Process
The discrepancy process term $\bm{\delta}_{t}$ is important because it respects the fact that our latent dynamical model represented by the process ${\bm{v}}_{NN,t}$ surely does not completely account for all of the latent spatio-temporal dependence in real-world data (i.e., “all models are wrong”). This term could be accommodated by another NN, say $\bm{\delta}_{t}=\bm{\delta}_{NN,t}(\cal{S};\bm{\theta}_{\delta})$ as in zou2024correcting. Alternatively, it could be a time-varying stochastic process, e.g., $\bm{\delta}_{t}=f_{\delta}(\bm{\delta}_{t-1};\bm{\theta}_{\delta})$ such as the residual autoregresssive process to accommodate small scale dynamics in wikle2001spatiotemporal, or simply a time-independent spatial GP. More generally, this term could be represented by any spatio-temporal random effect distribution, say $\bm{\delta}_{t}\sim[\bm{\theta}_{\delta}]$.
@@ 5.3 Parameter Models
Depending on the models selected in the data and process stages, one can have a variety of parameters that require prior distributions, e.g., $\bm{\beta}_{t}|\bm{\theta}_{\beta}][\bm{\theta}_{w}][\bm{\theta}_{\delta}]$, etc. Typically, we require regularization for these parameters as there are likely to be quite a lot of them, or as mentioned in the context of physical-statistical models, we might have mechanistic information for processes (e.g., parameters must obey a turbulent scaling relationship as in wikle2001spatiotemporal). In general, many of the prior specifications that have been used for mechanistically motivated BHM models (for some examples, see wikle2003hierarchical; pagendam2014assimilating; zammit2014resolving) carry over to BHM-PINNs because the BHM-PINN takes the same form at the parameter level. In the B-PINN context, HBPINN consider a hierarchical structure on parameters for a gear-grinding physics model in which correlation is allowed between parameters (we note, however, that this hierarchical framework is solely for the parameters of the physics model, distinct from the BHM-PINN approach).
Deep BHMs as described here have multiple hyperparameters. These correspond to specifications of the parameters for the priors in the previous stages as well as model architecture (e.g., number of hidden units, levels, etc.) in the case of B-PINNs. Depending on the application, these hyperparameters can be given vague proper prior distributions or are fixed (i.e., point mass priors) and the sensitivity to the choice is examined.
@@ 5.4 Implementation Considerations
As with many mathematical and statistical models, model identifiability must be be addressed when implementing a BHM-PINN, specifically with regard to the NN and discrepancy terms. For example, in (13), if $\bm{\delta}_{t}$ is parameterized using a GP, there is potential for confounding between this term and the PINN term, ${\bm{v}}_{NN,t}$. One way to mitigate this issue is to restrict the spatial (temporal, or spatio-temporal) range of the GP to operate on a different scale than the dynamics imposed on $\bm{v}_{NN,t}$. This can be accomplished by an informative prior distribution on the argument(s) of the kernel function for the GP.
@@
All the implementation challenges for B-PINNs are present for BHM-PINNs. For example, it is not clear how to select the NN architecture in general, but see toscano2025pinns for a review of various NN architectures that have been considered in the context of PINNs. Simple NNs, such as narrow 3-layer feed-forward NNs, are often used to fit B-PINNs. Not only is this more computationally tractable than a more complex model, it is also related to the bias-variance trade-off, where wide and deep NNs can be over-parameterized and have the potential to overfit. Empirical results often suggest narrow and shallow feed-forward NNs result in the best performance, although this may not be the case for complex dynamics.
@@
The primary computational tool that has been used for posterior inference in Bayesian hierarchical models is MCMC, such as Gibbs sampling (and variants like Metropolis-within-Gibbs sampling) (brooks2011handbook). Modern gradient-based methods, such as Hamiltonian Monte Carlo (HMC; neal2011mcmc) and extensions like the No-U-Turn Sampler (10.5555/2627435.2638586), have become prevalent in general for Bayesian inference. In the B-PINN literature (yang2021b), the two primary computational tools are HMC and variational inference (blei2017variational). In the case of variational Bayes, the mean-field approximation (i.e., product of independent normals) for posterior neural network weights might not be realistic in the sense that correlations (which are likely to exist in neural network parameters) are neglected, although variational GPs (tran2015variational) can help alleviate this issue. yang2021b found gradient-based MCMC methods often result in the best performance for B-PINNs and for our example presented below in Section 6, we use this approach.
In the context of the discrepancy term, we note that GP models can be computationally difficult due to cubic complexity of inversion and determinant calculations. This is a well-researched topic in spatial statistics and there are several ways to mitigate this issue (e.g., see the review in heaton2019case).
6 Simulation Experiment: BHM-PINN
@@
We consider a simulation experiment to illustrate a specific BHM-PINN. In particular, we consider a latent mechanistic dynamic process described by a one-dimensional (in space) Burgers’ equation (bateman1915some; burgers1948mathematical), which is a classic nonlinear PDE given by
$\frac{\partial{v}(s,t)}{\partial t}+{v}(s,t)\frac{\partial{v}(s,t)}{\partial s}-\lambda\frac{\partial^{2}{v}(s,t)}{\partial s^{2}}=0,$ (14)
@@
where $\lambda$ is the diffusion coefficient. The Burgers’ equation is often used to illustrate dynamics because it is a very simple nonlinear advection-diffusion representation of a fluid dynamics problem that can model features like nonlinear wave propagation and shock formation while remaining relatively easy to analyze and solve. Here, we also include a fixed-effects mean term in the process representation, along with a Gaussian process discrepancy to reflect a situation where the true process is not just a known mechanistic model. We then assume this process corresponds to the latent mean of a Poisson count data generating model. The details of the simulation are given below.
@@
Figure 1: Simulated latent process $u(s,t)$ (left) and components over one-dimensional space and time: Burgers’ equation v(s,t){v}(s,t) (second from left), mean process $\mu(s,t)$ (second from right), and discrepancy process $\delta(s,t)$ (right).
@@
    Process Mean ($\mu(s,t)$): We consider two covariates in the mean process, $\bm{x}(s,t)=(x_{1}(s,t),x_{2}(s,t))^{\top}$, which are simulated from a zero-mean spatio-temporal GP with exponential covariance function $k_{x}((s,t),(s^{\prime},t^{\prime}))=0.1\cdot\exp\bigg(-\frac{\sqrt{(s-s^{\prime})^{2}+(t-t^{\prime})^{2}}}{\ell_{s}}\bigg)$ where $\ell_{s}$ is set to $1/20$ the maximum observed spatio-temporal distance across all location-time pairs. Specifically, we choose $\ell_{s}=8.03/20$ and let $\bm{\beta}=(0.6,-0.4)^{\top}$. This is shown in the third panel of Figure 1.
@@
    Latent Dynamic Process ($v(s,t)$): For each spatial and temporal location, we generated the PDE solution to Burgers’ equation (14) numerically (with a spectral solver), denoted as v(s,t){v}(s,t). This numerical solution used initial condition ${v}(s,0)=\exp(-s^{2})$ for $s\in(-\pi,\pi)$ and 0 otherwise, with boundary conditions ${v}(-\pi,t)={v}(\pi,t)=0$. This is shown in the second panel of Figure 1.
@@
    Discrepancy Process ($\delta(s,t)$): For each time point a spatial discrepancy field was drawn as an independent, zero-mean Gaussian process, i.e., $\delta(s,t)\sim\text{GP}({0},k_{\delta}(s,s^{\prime}))$, where $k_{\delta}(s,s^{\prime})=\sigma^{2}_{\delta}\exp\bigg(-\frac{(s-s^{\prime})^{2}}{2\ell_{\delta}^{2}}\bigg)$ is a squared-exponential covariance function. Specifically, we select $\sigma^{2}_{\delta}=0.05$, and $\ell_{\delta}=0.15$. This is shown in the fourth panel of Figure 1.
@@
    Simulated Process ($u(s,t)$): The simulated process at each spatial and temporal coordinate is then given as the sum of the mean, latent dynamic process, and discrepancy: $u(s,t):=\bm{x}(s,t)^{\top}\bm{\beta}+{v}(s,t)+\delta(s,t)$. This is shown in the fourth panel of Figure 1.
@@
    Simulated Observations ($z(s,t)$): We then simulated observations from a random Poisson distribution, $z(s,t)\sim\mbox{Poisson}(\textrm{rate}=\exp(u(s,t)))$ (Figure 2 left panel). We assumed that observations at 30% of the space coordinates are missing at random for all times (Figure 2 right panel). Note, data missing over space for all times is a realistic scenario for simulating environmental processes with incomplete observation networks and can be more challenging to model than cases where data are missing completely at random in space and time.
@@
We consider the following BHM to fit these simulated data. First, the data, process, and discrepancy models are given by:
$\text{Data Model}:$ $\quad\bm{z}_{t}\sim\mbox{Poisson}(\exp(\bm{H}_{t}\bm{u}_{t}))$
$\text{True Process}:$ $\quad u(s,t)=\bm{x}(s,t)^{\top}\bm{\beta}+{v}_{NN}(s,t)+\delta(s,t)$
@@
$\text{Latent dynamics ($v_{NN}(s,t;\bm{\theta}_{W})$}:$ $\quad\frac{\partial{v}_{NN}(s,t)}{\partial t}\sim\text{N}(v_{NN}(s,t)\frac{\partial{v}_{NN}(s,t)}{\partial s}-\lambda\frac{\partial^{2}{v}_{NN}(s,t)}{\partial s^{2}},\sigma^{2}_{r})$
$\quad{v}_{NN}(s,0)\sim\text{N}(\exp(-s^{2}),\sigma^{2}_{ic})\quad\text{(Initial condition)}$
$\quad{v}_{NN}(-\pi,t),{v}_{NN}(\pi,t)\sim\text{N}(0,\sigma^{2}_{bc})\quad\text{(Boundary condition)}$
$\text{Discrepancy process}:$ $\quad\delta(s,t)\sim\text{GP}(0,k_{\delta}(s,s^{\prime}))$
$\quad k_{\delta}(s,s^{\prime})=\sigma^{2}_{\delta}\exp\bigg(-\frac{(s-s^{\prime})^{2}}{2\ell_{\delta}^{2}}\bigg).$
@@
In this case, $\bm{H}_{t}$ is an incidence matrix that accommodates missing observations (e.g., see the discussion in wikle2019). In addition, the following prior distributions are specified:
$\bm{\beta}$ $\sim\text{N}(\bm{\mu}_{\bm{\beta}},c_{\bm{\beta}}\bm{I}),\qquad$ $\lambda\sim\text{N}(\mu_{\lambda},\sigma^{2}_{\lambda})\cdot\text{I}(0\leq\lambda<\infty)$
$\ell_{\delta}$ $\sim\text{N}(\mu_{\ell},\sigma^{2}_{\ell})\cdot\text{I}(\ell_{\ell}\leq\ \ell_{\nu}<u_{\ell}),\qquad$ $\sigma^{2}_{\delta}\sim\text{N}(\mu_{\delta},\gamma^{2}_{\delta})\cdot\text{I}(\ell_{\delta}<\sigma^{2}_{\delta}<u_{\delta})$
$\bm{\theta}_{W}$ $\sim\text{N}(\bm{\mu}_{W},c_{W}\bm{I}),$